# CompAgent: An Agentic Framework for Visual Compliance Verification

### v1

@@ Appendix C Routing-based Architecture
Appendix D More Qualitative Results
[GT: Unsafe — Agent: Unsafe — LlavaGuard: Safe] GT:
@@ Appendix D More Qualitative Results
[GT: Unsafe — Agent: Unsafe — LlavaGuard: Safe] GT:
@@
[GT: Unsafe — Agent: Unsafe — LlavaGuard: Safe] GT:
@@
[GT: Safe — Agent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Safe — Agent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Safe — Agent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Safe — Agent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Safe — Agent: Unsafe — LlavaGuard: Safe] GT:
@@
[GT: Safe — Agent: Unsafe — LlavaGuard: Safe] GT:
@@
[GT: Safe — Agent: Unsafe — LlavaGuard: Safe] GT:
@@
[GT: Unsafe — Agent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Unsafe — Agent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Unsafe — Agent: Safe — LlavaGuard: Unsafe] GT:

### v2

@@ 8.4 Agentic Prompt
9 Routing Architecture
We illustrate our routing architecture in Fig. 4. A routing prompt drives Claude Sonnet 3.5 v2 to select and activate either category-based or policy-based nodes, which specialize in content-specific or policy-specific rules, respectively. Metadata produced by the activated nodes is then consolidated by a metadata fusion module. Finally, a multi-modal compliance analyzer (Claude Sonnet 3.5 v2) integrates the fused metadata with the compliance policy and input image to generate the final decision. In our experiments, we adopt the category-based routing configuration.
10 More Qualitative Results
[GT: Unsafe — CompAgent: Unsafe — LlavaGuard: Safe] GT:
@@ 10 More Qualitative Results
[GT: Unsafe — CompAgent: Unsafe — LlavaGuard: Safe] GT:
@@
[GT: Unsafe — CompAgent: Unsafe — LlavaGuard: Safe] GT:
@@
[GT: Safe — CompAgent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Safe — CompAgent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Safe — CompAgent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Safe — CompAgent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Safe — CompAgent: Unsafe — LlavaGuard: Safe] GT:
@@
[GT: Safe — CompAgent: Unsafe — LlavaGuard: Safe] GT:
@@
[GT: Safe — CompAgent: Unsafe — LlavaGuard: Safe] GT:
@@
[GT: Unsafe — CompAgent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Unsafe — CompAgent: Safe — LlavaGuard: Unsafe] GT:
@@
[GT: Unsafe — CompAgent: Safe — LlavaGuard: Unsafe] GT: